OpenStudy (madeleine2):

Tyler went to the supermarket to buy food for a food pantry. He has $36, and can carry up to 20 pounds of food in his backpack. Pasta costs $1 for a 1-pound package. Pasta sauce costs $3 for a 1.5 pound jar. Let x = the number of packages of pasta and y = the number of jars of pasta sauce. Identify each point as either a solution to the system or not a solution to the system of inequalities. (1, 12) (2, 10) (4, 5) (6, 8.25) (12, 8) (18, 6)

OpenStudy (bonnieisflash1.0):

sbcafk

OpenStudy (joshmorris):

p4tij

OpenStudy (vuriffy):

Hello.

OpenStudy (vuriffy):

(1, 12) (2, 10) (4, 5) (6, 8.25) (12, 8) (18, 6)

OpenStudy (vuriffy):

You have to test them for money limit and pound limit.

OpenStudy (vuriffy):

(1, 12) 1(1) + 3(12) = 1 + 36 = $37. Which is proven not to be a solution already for surpasses the money limit.

OpenStudy (madeleine2):

ok i plugged in the numbers like you did for the next one 1(2) + 3(10) but i dont know how to know if something is a solution or not @Vuriffy